OpenStudy (anonymous):

solve by the elimination method 0.3x - 0.2y=4 0.5x + 0.3y = 71/23

OpenStudy (amistre64):

yikes!! thats scary :)

OpenStudy (anonymous):

0.3x-0.2y=4 0.5x+0.3y=71/23 so, first you have to get rid of the y from both equations

OpenStudy (anonymous):

right! i'm clueless, the fraction throws me off!

OpenStudy (anonymous):

It is simple: multiply both equations, so that the y members of both of them are eliminated when added so in your case, we can multiply the first equation by 3, to get 0.9x-0.6y=12 and now we need to multiply the second equation in such a way that when coresponding members of both of them are added, y is eliminated (equals zero) so, we have to multiply it by 2 to get 1x+0.6y=142/23

OpenStudy (anonymous):

now, we have to add all coresponding members to eliminate y, so 0.9x-0.6y=12 + 1x+0.6y=142/23 _________________________ 1.9x+0=418/23 =1.9x=418/23

OpenStudy (anonymous):

ok, i follow you so far ....

OpenStudy (anonymous):

so, now lets solve for x 1.9x=418/23 23*1.9x=418 43.7x=418 x=418/43.7 x=9.56521

OpenStudy (anonymous):

i have to solve it as a fraction, so i'm plugging in 418/43.7 to solve for y ....

OpenStudy (anonymous):

now to get y we just take any equation, and substitute x into it, and solve for y 0.3x-0.2y=4 0.3*9.56521-0.2y=4 2.869565217-0.2y=4 -0.2y=4-2.869565217 -0.2y=1.130434 y=1.130434/-0.2 y=-5.65

OpenStudy (anonymous):

if I helped in anyway, I would apreciate if you would fan me (reload this page, and click a blue Become a fan button near my name)

OpenStudy (anonymous):

x~~9.56522, y~~-5.65217

OpenStudy (anonymous):

yes, thats correct

OpenStudy (anonymous):

what is y as a fraction?

OpenStudy (anonymous):

that is y as a decimal

OpenStudy (anonymous):

my answer has to be in the form of a fraction

OpenStudy (anonymous):

to get it as a fraction, just do same steps as I showed you, just keep pluging in fractions all the way from the beginning until now, and you can get a fraction

OpenStudy (anonymous):

x=220/23

OpenStudy (anonymous):

ok, thx! and i became a fan while you were working on my problem, thx!

OpenStudy (anonymous):

y=-130/23

OpenStudy (anonymous):

Got the fraction?

OpenStudy (anonymous):

x=220/23 y=-130/23

OpenStudy (anonymous):

great, thx!

OpenStudy (anonymous):

You are welcome, and look at the plot that will help you understand

OpenStudy (anonymous):

If you any more problems, you can send it my way, I will be happy to answer

OpenStudy (anonymous):

great, thx for your help!